BSW Corporation has a bond issue outstanding with an annual coupon rate of 5.2 percent paid quarterly and four years remaining until maturity. The par value of the bond is $1,000. Determine the fair present value of the bond if market conditions justify a 10.5 percent, compounded quarterly, required rate of return. (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16))

  

  Fair present value $   
Answer #1

Information provided:

Par value= future value= $1,000

Coupon rate= 5.2%/4= 1.30% per quarters

Coupon payment= 0.0130*1,000= $13

Time= 4 years*4= 16 quarters

Required rate of return= 10.5%/4= 2.6250% per quarter

The present value is computed by entering the below in a financial calculator:

FV= 1,000

PMT= 13

N= 16

I/Y= 2.6250

Press the CPT key and PV to calculate the present value:

The value obtained is 828.69.

Therefore, the fair present value is $828.69.

In case of any query, kindly comment on the solution